Calculate Thc pk of a osu o solution
Answer #1

Hi,

The given species can be dissociated as

NaClO >> Na+ + ClO-

Now,
ClO- + H2O <----> HClO + OH-
The constant of this reaction also called hydrolysis constant is Kw/ Ka = 1.x 10^-14 / 3.2 x10^-8 = 3.1 x 10^-7
3.1 x 10^-7 = x^2 / 0.5 -x
solving above expression for x
x = [OH-] = 4.29*10-4 M
log(OH) = pOH

pOH = 3.36
pH = 14 - 3.36 = 10.63

Thank You